If profit of 40% on selling price is equivalent to what percent profit on cost price.
Answer by RAY100(1637)

Let sp = c + p,,, selling price = cost + profit
.
given,,,p/sp = 40 % =.40
.
subst, p = sp -c
.
(sp-c) / sp = .4
.
sp - c = .4 sp
.
.6 sp = c
.
c /sp = .6 = 60 %
.
check, c/sp + p/sp =1, or c + p =sp ,,,,,ok